# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ 2.3.3 Thin radiative MADs
For even thinner MADs, we use a two-temperature (2T) radiation transport (GRRMHD) simulation of Liska et al. (2022) that results in a disk with a thermal scale height of $h/r\approx 0.03$. This simulation uses the M1 radiation scheme implemented in H-AMR, including Compton cooling (Liska et al., 2022). To capture the inner regions close to the BH and resolve the thin disk structure, the simulation uses adaptive mesh refinement (AMR) on a base grid of $1020\times 432\times 288$ cells. This resolution is then quadrupled at $5r_{\rm g}\leq r\leq 120r_{\rm g}$ and $|\theta-\pi/2|\leq 0.13$ to achieve an effective resolution of $4080\times 1728\times 1152$ cells in the disk proper. The BH spin is $a=0.9375$. We refer to this simulation as model Ra0.94. To improve our BH spin sampling, we restart this model, Ra0.94, at $t=188839r_{\rm g}/c$ for two different spin values, $a=0.3$ (Ra0.3) and $a=0.4$ (Ra0.4). This is the same as the start time of our analysis of Ra0.94 (see Table 1). We refer to these simulations collectively as $h/r=0.03$ $2T$. Table 1 summarizes the resolutions and scale heights.
2.4 Calculating fluxes
@@
$T^{\mu\nu}=T^{\mu\nu}_{\rm HD}+T^{\mu\nu}_{\rm EM},$ (12)
where the hydrodynamic part is given by
$T^{\mu\nu}_{\rm HD}=(\rho+u_{\rm g}+p_{\rm g})u^{\mu}u^{\nu}+pg^{\mu\nu},$ (13)
where $\rho$ is the gas density, $u$ is the internal energy, $p$ is the pressure, all measured in the fluid frame, $u^{\mu}$ is the contravariant 4-velocity, and $g^{\mu\nu}$ is the contravariant metric. The electromagnetic part of the stress-energy tensor is given by
$T^{\mu\nu}_{\rm EM}=b^{2}u^{\mu}u^{\nu}+\frac{1}{2}b^{2}g^{\mu\nu}-b^{\mu}b^{\nu},$ (14)
where $b^{\mu}$ is the magnetic field $4$-vector.

### v2

@@ II.3.3 Thin radiative MADs
To model radiatively cooled MADs with realistic thermodynamics, we use a two-temperature (2T) radiation transport (GRRMHD) simulation of [51] that results in a disk with a thermal scale height of $h/r\approx 0.03$. This simulation uses the M1 radiation scheme implemented in H-AMR, including the Compton cooling [51]. To resolve the thin disk structure, the simulation uses a static mesh refinement (SMR) on a base grid of $1020\times 432\times 288$ cells. We then quadruple this resolution at $5r_{\mathrm{g}}\leq r\leq 120r_{\mathrm{g}}$ and $|\theta-\pi/2|\leq 0.13$ to achieve an effective resolution of $4080\times 1728\times 1152$ cells in the disk proper. The BH spin is $a=0.9375$. We refer to this simulation as model Ra0.94. To improve our BH spin sampling, we restart this model, Ra0.94, at $t=188839r_{\mathrm{g}}/c$ for two different spin values, $a=0.3$ (Ra0.3) and $a=0.4$ (Ra0.4). This is the same as the start time of our analysis of Ra0.94 (see Table 1). We refer to these simulations collectively as $h/r=0.03$ $2T$. Table 1 summarizes the resolutions and scale heights.
II.4 Calculating fluxes
@@
$T^{\mu\nu}=T^{\mu\nu}_{\mathrm{HD}}+T^{\mu\nu}_{\mathrm{EM}},$ (12)
where the hydrodynamic (HD) part is
$T^{\mu\nu}_{\mathrm{HD}}=(\rho+u_{\mathrm{g}}+p_{\mathrm{g}})u^{\mu}u^{\nu}+p_{\mathrm{g}}g^{\mu\nu}.$ (13)
Here, $\rho$ is gas density, $u_{\mathrm{g}}$ is the internal energy, and $p_{\mathrm{g}}$ is the pressure of the gas, all measured in the fluid frame, $u^{\mu}$ is the contravariant 4-velocity, and $g^{\mu\nu}$ is the contravariant metric. The electromagnetic (EM) part of the stress-energy tensor is given by
$T^{\mu\nu}_{\mathrm{EM}}=b^{2}u^{\mu}u^{\nu}+\frac{1}{2}b^{2}g^{\mu\nu}-b^{\mu}b^{\nu},$ (14)
where $b^{\mu}$ is the magnetic field four-vector and $b^{2}=b^{\mu}b_{\mu}=2p_{\mathrm{mag}}$ is twice the magnetic pressure.